OpenStudy (pottersheep):

f(x) = x^2. How would I find the coordinates of the graph f(x-3) +1 ?

OpenStudy (pottersheep):

Do i substitute x^2 into f(x-3) +1 whenever x appears ? so 3(x^2 -3 ) + 1????

OpenStudy (saifoo.khan):

f(x) = x^2 f(x-3)+1 = (x-3)^2 +1 = x^2 +9+1

OpenStudy (anonymous):

just one modification to it 3(x-3)^2 +1

OpenStudy (saifoo.khan):

from where do u get that 3

OpenStudy (pottersheep):

Thanks

OpenStudy (anonymous):

Your question asks to identify the coordinates of the graph; the graph is going to be an upward parabola, which has been translated to the right 3 units and up 1 unit. i.e., the vertex is (3,1) and is constructed as per the usual conventional upward turning parabola.